Finance Statement

The financial statements for MHM Bank (MHM) are shown below:
 

Balance Sheet MHM Bank
AssetsLiabilities and Equity
Cash and due from banks$2,050
Demand deposits$10,750
Demand deposits at other FIs
2,400
Small time deposits
10,480
Investments
6,340
Jumbo CDs
7,800
Federal funds sold
3,120
Federal funds purchased
600
Loans (less reserve for loan losses of 3,700)
20,320
Other liabilities
3,580
Premises
2,400
Equity
3,420
Total assets$36,630
Total liabilities/equity$36,630

 

Income Statement MHM Bank
Interest income$4,318
Interest expense
2,284
Provision for loan losses
1,400
Noninterest income
830
Noninterest expense
1,105
Taxes
365


a. Calculate the dollar value of MHM’s earning assets.
b. Calculate the dollar value of MHM’s interest-bearing liabilities.
c. Calculate MHM’s spread. (Do not round intermediate calculations. Round your answer to 3 decimal places. (e.g., 32.161))
d. Calculate MHM’s interest expense ratio. (Round your answer to 3 decimal places. (e.g., 32.161))
Answer #1

Requirement-(a):

$$ \begin{array}{lr} \text { Investments } & \$ 6,340 \\ \text { Federal funds sold } & \$ 3,120 \\ \text { Loans (less reserve for loan losses of } 3,700) & \$ 20,320 \\ \text { Eearning Assets } & \$ 29,780 \\ \hline \hline \end{array} $$

Requirement-(b):

$$ \begin{array}{lr} \text { Small Time Deposits } & \$ 10,480 \\ \text { Jumbo CDs } & \$ 7,800 \\ \text { Federal Funds Purchased } & \$ 600 \\ \hline \text { Interest bearing liabilities } & \$ 18,880 \\ \hline \hline \end{array} $$

Requirement-(c):

$$ \begin{array}{lcll} \text { Eearning Assets } & \$ 29,780 & & \text { [Ref: Requirement-(a)] } \\ \text { Interest income } & \$ 4,318 & & \\ \text { Interest Income Spread } & & 14.500 \% & & \text { [Interest income / Earning assets }=\$ 4,318 / \$ 29,780=0.145=14.500 \%] \\ & & & \\ \text { Interest bearing liabilities } & \$ 18,880 & & \text { [Ref: Requirement-(b)] } \\ \text { Interest Expense } & \$ 2,284 & & \\ \text { Interest Expense Spread } & & 12.097 \% & \text { [Interest expense / Interest bearing liabilitites = } \$ 2,284 / \$ 18,880=0.12097=12.097 \%] \\ \text { MHM Bank's SPREAD } & & 2.402 \% & \text { [Income Spread - Expense Spread = } 14.500 \%-12.097 \%=2.402 \%] \end{array} $$

Requirement-(d):

$$ \begin{array}{lr} \text { Interest income } & \$ 4,318 \\ \text { Non Interest income } & \$ 830 \\ \hline \text { Total Income } & \$ 5,148 \\ \text { Interest Expense } & \$ 2,284 \\ \hline \text { Interest Expense Ratio } & 44.367 \% \\ \hline \hline \end{array} $$

answered by: Crispan